# FGBERT: Function-Driven Pre-trained Gene Language Model for Metagenomics

### v1

@@ 3.2 Triplet Enhanced Metagenomic Contrastive Learning (TEM-CL)
Although MGM can learn the gene-protein contextual relationship within metagenomic data, we find its ability to identify sequences with same function to be lacking, which is common in Many-to-One scenarios. During the Enzyme Commission (EC) number annotation of partial metagenomic sequences, we observe that sequences with the same EC number are close to each other in feature space and, vice versa, farther apart. This finding motivates us to adopt a contrastive learning approach to capture the functional relationships between gene classes, enabling different genes with similar functions to cluster together and further optimize model training.
@@
The common contrastive learning objective is to learn an embedding function $\mathcal{F}$ such that the distance between positive pairs is less than that of negative pairs: $d(\mathcal{F}(x_{a}),\mathcal{F}(x_{p}))<d(\mathcal{F}(x_{a}),\mathcal{F}(x_{n}))$, where $d(\cdot,\cdot)$ is a distance function (e.g., Euclidean distance) defined on the embedding space. We adopt SupCon-Hard loss [29] to thus consider multiple positive and negative samples for each anchor, which encourages the model to mine from difficult samples and enhances the robustness and generalization of the model. Additionally, the data augmentation and negative sampling modules are also included to create positive and negative samples and improve the model’s capacity to recognize commonalities among gene classes.
@@
Triplet Sampling How to sample the triplets is crucial to learning a well-organized embedding space. For each gene group $\mathcal{G}$, as an anchor gene $x_{i}$ within a gene batch $I$, a mutation strategy is proposed to augment orphan sequences (i.e., functions associated with individual sequences) to generate a large number of pairs of positive samples $x_{j(i)}\in\mathcal{G}_{i}^{+}$ where $\mathcal{G}_{i}^{+}$ is the positive samples set of the anchor $x_{i}$. Concretely, 10 random mutations are performed on each gene sequence, with mutation ratios randomly generated according to a standard normal distribution, and the number of mutations is calculated based on the length of the sequence. This process aims to generate new sequences that are functionally similar to the original sequence but sequentially different, thus providing additional training data for the model and improving the predictive power and accuracy of orphan EC numbers.
Hard Negatives Sampling Previous studies [19] have shown that another key to successful contrastive learning is the balance between the triviality and the hardness of the sampled triplets. For the negative sample pair $x_{k(i)}\in\mathcal{G}_{i}^{-}$ where $\mathcal{G}_{i}^{-}$ is the negative samples set of the anchor $x_{i}$.
To begin, we calculate cluster centers for each EC number sequence by averaging the embedded representations of all sequences within the same functional group. These cluster centers represent the centroids of each functional group. Subsequently, we compute Euclidean distances $d(\cdot)$ based on these centers. When selecting negative samples, we choose samples that are similar to the anchor in latent space but belong to different clusters, rather than randomly selecting them, to increase the challenge of learning. The triplet loss $\mathcal{L}_{\text{Tri}}(x_{i},\{x_{j(i)}\}_{j=1}^{N_{j}},\{x_{k(i)}\}_{k=1}^{N_{k}})$ is defined as follows:
$\mathcal{L}_{\text{Tri}}=-\textstyle\sum\limits_{i\in I}\log(\frac{1}{|\mathcal{G}_{i}^{+}|}\textstyle\sum\limits_{j\in\mathcal{G}_{i}^{+}}\frac{\exp(S_{z_{i},z_{j(i)}}/\tau)}{\mathcal{G}_{i}})$ (3)
For all negative samples in $\mathcal{G}_{i}^{-}$, we calculate the weight $w_{x_{k}}=\frac{1}{d(x_{i},x_{k})}$ for each $x_{k}\in\mathcal{G}_{i}^{-}$ based on its distance relative to anchor $x_{i}$. We then normalize these weights to obtain the probability $p_{x_{k}}$: $p_{x_{k}}=\frac{w_{x_{k}}}{\textstyle\sum_{x_{m}\in\mathcal{G}_{i}^{-}}w_{x_{m}}}$.
$\mathcal{G}_{i}=\textstyle\sum\limits_{x_{j(i)}}\exp(S_{z_{i},z_{j(i)}}/\tau)+\textstyle\sum\limits_{x_{k(i)}}p_{x_{k}}\exp(S_{z_{i},z_{k(i)}}/\tau)$ (4)
@@
where $\tau$ is the temperature hyper-parameter and $S$ is the similarity function, typically defined by cosine similarity.
@@
Finally, MGM and TEM-CL constitute a unified pre-training framework with a total loss:
$\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{MGM}}+\lambda\mathcal{L}_{\text{Tri}}$ (5)
@@ 4 Experiments
We pre-train FGBERT on a large amount of metagenomic data and comprehensively assess its generalization on different datasets, from thousands to hundreds of thousands, which are described in detail in Table 1. Our goal is to validate the performance of our proposed model on four levels of tasks: (1) Genomic Structure Analysis; (2) Functional Gene Prediction (3) Pathogenicity Potential Assessment and (4) Nitrogen Cycle Prediction.
4.1 Pre-training Datasets
We use MGnify updated 2023_02, containing 2,973,257,435 protein sequences. MGnify [50] is a microbial genome dataset that contains genome sequences from different microbial communities. Detailed dataset information is shown in Appendix 5.
Table 1: Description of Experimental Datasets. #Seq. and #Class means the number of sequences and categories in each dataset.
Dataset
Description
#Seq.
#Class
E-K12
Gene Operons
4,315
1,379
CARD-A
AMR Gene Family
1,966
269
CARD-D
Drug Class
1,966
37
CARD-R
Resistance Mechanism
1,966
7
VFDB
Virulence Factors
8,945
15
ENZYME
Enzyme Functions
5,761
7
PATRIC
Pathogenic Genes
5,000
110
NCycDB
Nitrogen Cycling Genes
213,501
68
[…]
Method Operons ARG Prediction Virus Enzyme Pathogen N-Cycle
E-K12 CARD-A CARD-D CARD-R VFDB ENZYME PATRIC NCycDB
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
RF
20.2
34.8
22.4
35.3
36.1
49.0
47.8
57.6
22.4
38.5
33.6
41.2
25.3
29.8
67.0
71.7
SVM
38.6
45.2
27.6
40.5
33.6
47.2
43.3
66.2
28.0
41.4
31.3
43.6
26.6
31.2
66.9
70.3
KNN
39.9
41.0
36.9
54.4
36.4
51.3
36.2
63.5
27.3
47.1
31.4
42.9
11.0
27.4
68.8
73.2
LSTM
40.4
42.5
47.1
60.3
39.1
62.3
47.5
84.2
36.7
66.3
42.8
51.0
41.3
49.7
71.9
81.2
BiLSTM
38.2
43.8
47.4
61.9
43.5
58.1
58.9
80.3
46.1
72.1
38.7
50.2
43.3
48.5
82.0
88.4
VT
43.3
47.8
57.1
70.0
49.8
68.1
55.7
86.4
58.0
81.0
68.2
75.8
49.8
57.3
84.5
90.7
HyenaDNA
42.4
47.1
50.9
68.2
53.6
78.1
66.2
88.1
61.0
70.4
79.6
83.6
51.1
57.6
92.4
96.0
ESM-2
38.2
42.5
57.2
71.4
56.0
82.1
68.2
90.0
60.7
84.4
92.5
96.7
56.0
67.5
95.8
96.1
NT
45.1
44.8
58.5
72.0
56.2
80.2
68.0
90.3
58.3
71.6
74.1
76.7
46.1
61.9
75.1
86.5
DNABERT2
51.7
52.4
65.2
79.8
51.5
78.7
61.2
88.6
58.2
82.3
85.4
85.2
52.9
60.6
88.6
95.7
Ours 61.8 65.4 78.6 90.1 57.4 85.2 69.4 91.4 75.7 90.2 99.1 98.8 99.3 99.0 99.5 99.2
4.2 Genomic Structure Analysis
Gene Operons Prediction is to identify the transcription factor binding sites that have the strongest correlation with operon regulation in the gene regulatory network, which helps us to understand the mechanism and network of gene regulation.
Dataset. The transcription factor binding sites with the strongest correlation to the operon in E.coil K12 RegulonDB dataset (E-K12) [51] consists of 4315 operons. Detailed dataset information is listed in Appendix 6.
Results. Table 2 shows that our model achieves the SOTA results for the operon prediction task. From the attention heatmap in Figure 4, it can be observed that gene operon tolC has high attention weight with the operons ygiB and ygiC and yqiA. This suggests a significant interaction among these operons, indicating the presence of a shared genetic operon tolC-ygib. Biological research supports that these operons are indeed related to the DUF1190 domain-containing protein YgiB [28].
Figure 4: Visualization of the attention heatmap from the first head of the last (19th) attention layer. Darker blue color indicates higher attention weight.
4.3 Functional Gene Prediction
Antimicrobial Resistance Genes (ARG) Prediction is crucial for understanding antimicrobial resistance genes and facilitates the identification of resistance mechanisms. However, existing methods suffer from high false-positive rates and specific category bias [26, 5]. Therefore, there is an urgent need for deep learning methods to rapidly and accurately predict the presence of ARG in metagenomic data.
Dataset. CARD [25] is a comprehensive ARG database in which each gene is associated with a specific AMR Gene Family (CARD-A: 269 categories in total), Drug Class (CARD-D: 37 categories in total), and Resistance Mechanism(CARD-R: 7 categories in total). Therefore, our model performs a three-category classification prediction task for each gene sequence in CARD, respectively.
Results. The prediction on CARD-A category is especially noteworthy in Table 2. Because this category covers up to 269 different classifications, resulting in a long-tail issue. This requires the model to understand the biological properties of gene sequences and accurately annotate the sequences [40]. To address the long-tail problem, we adjust the data sampling strategy to increase the sampling frequency of fewer samples to improve the model’s prediction accuracy. Table 3 demonstrates the prediction accuracy of FGBERT on CARD-R category. This result shows that our model has better classification results for both majority and minority classes with more than 85% accuracy. Appendix 9 shows that FGBERT has better classification performance for all 269 categories of AMR Gene Family, with 100% data classification accuracy for majority categories such as CTX, ADC, CMY, etc. and minority categories such as AXC, CRH, KLUC, and so on.
Table 3: Detailed Classification Results of FGBERT Model on Resistance Mechanism Category Prediction, including the total number of categories, the number of correct predictions, and the accuracy rate for each class.
Class
Name
Total
Count
Correct
Num
Correct
Ratio
@@ 4.3 Functional Gene Prediction
Virulence Factors (VF) Prediction is to detect microbial elements like bacterial toxins, which enhance pathogen infectivity and exacerbate antimicrobial resistance. Existing methods for analyzing metagenomic analyses are insufficient, especially tools for co-predicting ARGs and VFs are scarce, and also suffer from threshold sensitivity [61].
Dataset. VFDB [8] dataset includes the major VFs of the most characterized bacterial pathogens, as well as the structural features, functions, and mechanisms of these pathogens. We use VFDB core dataset, including 8945 VF sequences and 15 VF categories.
Results. According to the results reported in Table 2, our model achieves the SOTA results on VFDB. Compared to the genomic pre-trained model, our advantage is relatively obvious. For example, M.F1 improves by 30% and W.F1 improves by 9.5% compared to DNABERT2. This further highlights the challenge of directly transferring the genome pre-trained model to the metagenomic domain for precise functional annotation. Conversely, ESM-2, as a PLM, outperformed the genome pre-trained model, emphasizing the significance of harnessing the inherent protein information within metagenomic data.
Enzyme Function Prediction is critical for understanding metabolism and disease mechanisms in organisms. While traditional methods rely on time-consuming and labor-intensive biochemical experiments, advanced technologies can be used to efficiently and accurately predict large-scale genomic data.
Dataset. ENZYME [7] is a repository of information related to the nomenclature of enzymes. We organize 5761 data, 7 categories of ENZYME core dataset, each enzyme has a unique EC number [44].
@@
Results. Our experimental results demonstrate FGBERT’s superior performance on ENZYME dataset. It outperforms ESM-2, the second-highest method, by approximately 6.62% in M.F1 and 2.09% in W.F1, demonstrating its ability to discern distinct enzyme function characteristics. This observation highlights that our model not only captures gene-protein contextual relationships but also effectively models the relationships between sequences and functions within metagenomic data.
4.4 Pathogenicity Potential Assessment
Genome Pathogens Prediction is to assess the pathogenic potential of pathogens to cope with the public health risks caused by newly emerging pathogens. Hence, the development of accurate deep-learning algorithms for the precise identification of pathogens can improve the ability to respond to the threat of drug resistance.
Dataset. We use PATRIC core dataset [15], which has 5000 pathogenic bacterial sequences and 110 pathogenic bacterial classes.
Results. Table 2 displays the classification results for pathogenic bacteria species on PATRIC. The result shows that FGBERT can recognize key features of genera, resulting in better classification results than baselines. PATRIC dataset, which consists of a large number of categories with relatively fewer data, poses a challenge for models. Due to this complexity, baselines underperform, as they typically require more data to discern the subtle differences between numerous categories. In contrast, FGBERT stands out with M.F1 and W.F1 scores of 99.27% and 99.03% respectively, and appears to be robust against such data scarcity and category abundance, indicating an advanced learning capability suited for high-dimensional classification tasks. This observation highlights the advantage of leveraging protein-based gene representations for improved accuracy in functional annotation tasks.
4.5 Nitrogen Cycle Prediction
Nitrogen Cycling Process Prediction is to quantitatively study the functional genes related to the N cycle and linking them to environmental and ecological processes is one of the important focuses in environmental genomics research.
Dataset. NCycDB [55] contains a total of 68 genes (sub)families and covers eight N cycle processes with 213501 representative sequences at 100% identity cutoffs, each of which involves a specific gene family and is critical for environmental and ecosystem nitrogen homeostasis.
Results. Table 2 presents the classification results of FGBERT on NCycDB. The findings suggest that FGBERT can recognize important features of a particular N cycle and improve gene family classification by recognizing domains. On NCycDB, which consists of a substantial 210,000 data from 68 categories, we observe that the baselines perform comparatively better than on PATRIC. This is due to the larger data size per category, which allows baselines to better discriminate between the diverse categories, while FGBERT still outperforms with remarkable M.F1 and W.F1 scores of 99.49% and 99.22%, respectively. However, the pre-trained baselines take more time and larger memory for the tokenization process for large datasets, which is analyzed in Sec 4.7.

### v2

@@ 3.2 Triplet Metagenomic Contrastive Framework (TMC) for Many-to-One problem
Contrastive Learning. The MGM component enhances the model’s ability to learn contextual relationships between genes, helping to alleviate the OTM problem present in metagenomic data. However, when faced with the MTO problem, the model’s ability to describe the relationships between sequences with the same function remains weak. For instance, when annotating Enzyme Commission (EC) numbers for partial metagenomic sequences, we observe that sequences with the same EC number are close to each other in feature space, while sequences with different EC numbers are farther apart [29]. Therefore, we introduce a contrastive learning technique to capture the functional relationships between gene classes, enabling different genes with similar functions to cluster together and further optimize model training. Generally speaking, the objective of contrastive learning is to learn an embedding function $\mathcal{F}$ such that the distance between positive pairs is smaller than the distance between negative pairs:
$d(\mathcal{F}(x_{a}),\mathcal{F}(x_{p}))<d(\mathcal{F}(x_{a}),\mathcal{F}(x_{n})),$ (3)
@@
where $d(\cdot,\cdot)$ is a distance function (e.g., Euclidean distance) defined on the embedding space. We adopt SupCon-Hard loss [31] to consider multiple positive and negative samples for each anchor, encouraging the model to mine difficult samples and enhancing its robustness. Additionally, data augmentation and negative sampling modules are included to create positive and negative samples, further improving the model’s capacity to recognize commonalities among gene classes.
@@
Positives Sampling. The strategy for sampling triplets is crucial to learn a well-organized embedding space. For each gene group $\mathcal{G}$, as an anchor gene $x_{i}$ within a gene batch $I$, a mutation strategy is proposed to augment orphan sequences (i.e., functions associated with individual sequences) to generate a large number of pairs of positive samples $x_{j(i)}\in\mathcal{G}_{i}^{+}$, where $\mathcal{G}_{i}^{+}$ is the set of positive samples for anchor $x_{i}$. Specifically, 10 random mutations are performed for each gene sequence, with mutation ratios randomly generated according to a standard normal distribution. The number of mutations is calculated based on sequence lengths. This process aims to generate new sequences that are functionally similar to the original sequence but sequentially different, providing additional training data to improve the predictive power and accuracy of orphan EC numbers.
Hard Negatives Sampling Previous studies [19] have demonstrated that a critical component of successful contrastive learning is balancing the triviality and hardness of the sampled triplets. For the negative sample pair $x_{k(i)}\in\mathcal{G}_{i}^{-}$, where $\mathcal{G}_{i}^{-}$ represents the set of negative samples for the anchor $x_{i}$, this balance is particularly important. We determine the centers for each functional group by averaging the embeddings of all sequences within that group. Subsequently, we compute the Euclidean distances $d(\cdot)$ based on these centers. For negative sample selection, we choose samples that are similar to the anchor in latent space but from different clusters, thus increasing the learning difficulty compared to random selection. The triplet loss $\mathcal{L}_{\text{Tri}}(x_{i},\{x_{j(i)}\}_{j=1}^{N_{j}},\{x_{k(i)}\}_{k=1}^{N_{k}})$ is defined:
$\mathcal{L}_{\text{Tri}}=-\sum_{i\in I}\log({1}/{|\mathcal{G}_{i}^{+}|}\sum_{j\in\mathcal{G}_{i}^{+}}{\exp(S_{z_{i},z_{j(i)}}/\tau)}/{\mathcal{G}_{i}}).$ (4)
For all negative samples in $\mathcal{G}_{i}^{-}$, the probability of selecting each negative sample $x_{k}$ for the anchor $x_{i}$ as follows:
$\mathcal{G}_{i}=\sum_{x_{j(i)}}\exp(S_{z_{i},z_{j(i)}}/\tau)+\sum_{x_{k(i)}}p_{x_{k}}\exp(S_{z_{i},z_{k(i)}}/\tau),$ (5)
@@
where $\tau$ is the temperature hyper-parameter, and $S$ is the similarity function, typically cosine similarity. The term $p_{x_{k}}$ represents the probability of selecting the negative sample $x_{k}$ for the anchor $x_{i}$, calculated as $p_{x_{k}}={w_{x_{k}}}/{\sum_{x_{m}\in\mathcal{G}_{i}^{-}}w_{x_{m}}}$ with $w_{x_{k}}=\frac{1}{d(x_{i},x_{k})}$.
@@
Finally, MGM and TMC constitute a unified pre-training framework with a total loss:
$\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{MGM}}+\lambda\mathcal{L}_{\text{Tri}},$ (6)
@@ 4 Experiments
Table 2: Description of Experimental Datasets.
Task Dataset Description #Seq. #Class Gene Structure Prediction (Gene Level) E-K12 Gene Operons 4,315 1,379 Functional Prediction (Functional Level) CARD-A AMR Family 1,966 269 CARD-D Drug Class 1,966 37 CARD-R Resistance Mech. 1,966 7 VFDB Virulence Fact. 8,945 15 ENZYME Enzyme Func. 5,761 7 Pathogenicity Prediction (Bacteria Level) PATRIC Pathogenic Genes 5,000 110 Nitrogen Cycle Prediction (Environmental Level) NCycDB Cycling Genes 213,501 68
@@
We pre-train FGBERT on a large amount of metagenomic data and comprehensively assess its generalization on different datasets ranging from thousands to hundreds of thousands of sequences, as detailed in Tab. 2. Our model is tested across four task levels: (1) Gene Structure Analysis, (2) Functional Gene Prediction, (3) Pathogenicity Potential Assessment, and (4) Nitrogen Cycle Prediction. More detailed description of the downstream tasks can be found in Appendix Sec. C.4. We use the MGnify database (updated February 2023), which comprises 2,973,257,435 protein sequences from various microbial communities, detailed in Appendix Tab. 14.
4.1 Experiments Results on Four Level Downstream Tasks
Level 1 Task A: Gene Structure Analysis $\rightarrow$ Gene Operons Prediction. This task is to identify the transcription factor binding sites that are strongly correlated with operon regulation in the gene regulatory network, which helps us to understand the mechanism and network of gene regulation. The dataset used is the E. coli K12 RegulonDB dataset (E-K12) [55], which contains 4315 operons. Detailed information is listed in Appendix Tab. 15.
Table 3: Visualization of Attention.
Results Analysis. The attention heatmap in Fig. 3 shows that gene operon tolC has high attention weight with the operons ygiB and ygiC and yqiA. This suggests a significant interaction among these operons, indicating the presence of a shared genetic operon tolC-ygib. This inference finds support in biological research, which associates these operons with the DUF1190 domain-containing protein YgiB [30].
Level 2 Task B: Functional Gene Prediction $\rightarrow$ Antimicrobial Resistance Genes (ARG) Prediction. This task is crucial for understanding ARGs and facilitates the identification of resistance mechanisms. However, existing methods suffer from high false-positive rates and category bias [27, 5], necessitating the use of deep learning methods to rapidly and accurately detect ARG presence in metagenomic data. The CARD dataset [26] categorizes each gene into one of 269 AMR Gene Families (CARD-A), 37 Drug Classes (CARD-D), or 7 Resistance Mechanisms (CARD-R). Our model performs a three-category classification for each gene sequence in CARD, respectively.
Table 4: Classification Results on CARD-R.
ClassName Total Count Correct Num Correct Ratio
@@ 4.1 Experiments Results on Four Level Downstream Tasks
Results Analysis. FGBERT’s performance on CARD-A is significant, as shown in Tab. 5. This category’s broad range (269 classifications) creates a long-tail distribution, necessitating an understanding of the biological properties of gene sequences for accurate annotation [44]. To mitigate this issue, we adjust the data sampling strategy to increase the frequency of fewer samples, improving the model’s prediction accuracy. Tab. 4 demonstrates FGBERT’s high prediction accuracy for the CARD-R category, exhibiting superior classification results for both majority and minority classes, with over 85% accuracy. Appendix Tab. LABEL:tab:CARD_ours_AMR_ratio reveals FGBERT has better performance for all 269 AMR Gene Family categories, with 100% classification accuracy for majority categories like CTX, ADC, CMY, as well as for minor ones like AXC, CRH, KLUC.
[…]
Method Operons ARG Prediction Virus Enzyme Pathogen N-Cycle E-K12 CARD-A CARD-D CARD-R VFDB ENZYME PATRIC NCycDB M.F1 W.F1 M.F1 W.F1 M.F1 W.F1 M.F1 W.F1 M.F1 W.F1 M.F1 W.F1 M.F1 W.F1 M.F1 W.F1 RF 20.2 34.8 22.4 35.3 36.1 49.0 47.8 57.6 22.4 38.5 33.6 41.2 25.3 29.8 67.0 71.7 SVM 38.6 45.2 27.6 40.5 33.6 47.2 43.3 66.2 28.0 41.4 31.3 43.6 26.6 31.2 66.9 70.3 KNN 39.9 41.0 36.9 54.4 36.4 51.3 36.2 63.5 27.3 47.1 31.4 42.9 11.0 27.4 68.8 73.2 LSTM 40.4 42.5 47.1 60.3 39.1 62.3 47.5 84.2 36.7 66.3 42.8 51.0 41.3 49.7 71.9 81.2 BiLSTM 38.2 43.8 47.4 61.9 43.5 58.1 58.9 80.3 46.1 72.1 38.7 50.2 43.3 48.5 82.0 88.4 VT 43.3 47.8 57.1 70.0 49.8 68.1 55.7 86.4 58.0 81.0 68.2 75.8 49.8 57.3 84.5 90.7 HyenaDNA 42.4 47.1 50.9 68.2 53.6 78.1 66.2 88.1 61.0 70.4 79.6 83.6 51.1 57.6 92.4 96.0 ESM-2 38.2 42.5 57.2 71.4 56.0 82.1 68.2 90.0 60.7 84.4 92.5 96.7 56.0 67.5 95.8 96.1 NT 45.1 44.8 58.5 72.0 56.2 80.2 68.0 90.3 58.3 71.6 74.1 76.7 46.1 61.9 75.1 86.5 DNABERT2 51.7 52.4 65.2 79.8 51.5 78.7 61.2 88.6 58.2 82.3 85.4 85.2 52.9 60.6 88.6 95.7 Ours 61.8 65.4 78.6 90.1 57.4 85.2 69.4 91.4 75.7 90.2 99.1 98.8 99.3 99.0 99.5 99.2
Level 2 Task C: Functional Gene Prediction $\rightarrow$ Virulence Factors (VF) Prediction. This task is to detect microbial elements like bacterial toxins, which enhance pathogen infectivity and exacerbate antimicrobial resistance. Existing methods for metagenomic analysis, particularly those co-predicting ARGs and VFs, are inadequate and suffer from threshold sensitivity issues [65]. VFDB dataset [8] includes the major VFs of the most characterized bacterial pathogens, detailing their structural features, functions, and mechanisms. We use VFDB core dataset, including 8945 VF sequences and 15 VF categories. Results Analysis. Our model achieves the SOTA results on VFDB, as reported in Tab. 5. It significantly outperforms the genomic pre-trained model; for instance, M.F1 and W.F1 scores improve by 30% and 9.5%, respectively, compared to DNABERT2. This highlights the limitations of directly applying the genomic pre-trained model to metagenomic data for precise functional annotation. Conversely, ESM-2, as a PLM, excels by leveraging intrinsic protein information in metagenomic data, highlighting its effectiveness.
@@
Level 2 Task D: Functional Gene Prediction $\rightarrow$ Enzyme Function Prediction. This task is critical for understanding metabolism and disease mechanisms in organisms. While traditional methods rely on time-consuming and labor-intensive biochemical experiments, advanced technologies can offer efficient and accurate predictions for large-scale genomic data. ENZYME [7] is a repository of information related to the nomenclature of enzymes. We organize 5761 data, 7 categories of ENZYME core dataset, each enzyme has a unique EC number [49]. Results Analysis. Our experimental results demonstrate FGBERT’s superior performance on the ENZYME dataset. It outperforms ESM-2, the second-highest method, by approximately 6.62% in M.F1 and 2.09% in W.F1, demonstrating its ability to discern distinct enzyme function characteristics. This observation highlights that our model not only captures gene-protein contextual relationships but also effectively models the relationships between sequences and functions within metagenomic data.
Level 3 Task E: Pathogenicity Potential Assessment $\rightarrow$ Genome Pathogens Prediction. This task assesses the pathogenic potential of pathogens to cope with the public health risks caused by newly emerging pathogens. Accurate deep-learning algorithms are key for the precise identification of pathogens, improving the ability to respond to drug resistance threats. We use PATRIC core dataset [15], which has 5000 pathogenic bacterial sequences across 110 classes. Results Analysis. Tab. 5 shows FGBERT’s classification of pathogenic bacteria species within PATRIC, demonstrating superior performance over baselines by recognizing crucial genera features. The PATRIC dataset presents a significant challenge due to its large number of categories and sparse data. Baselines generally underperform because they require more data to discern the subtle differences between numerous categories. In contrast, FGBERT stands out with M.F1 and W.F1 scores of 99.27% and 99.03%, respectively. This robust performance indicates its advanced learning capability, making it well-suited for high-dimensional classification tasks and highlighting the benefits of using protein-based gene representations for enhanced functional annotation accuracy.
Level 4 Task F: Nitrogen Cycle Prediction $\rightarrow$ Nitrogen (N) Cycling Process Prediction. This task focuses on the functional genes related to the N cycle, linking them to environmental and ecological processes. NCycDB [59] contains 68 gene (sub)families and covers 8 N cycle processes with 213,501 representative sequences at 100% identity cutoffs, each involving a specific gene family. Results Analysis. Tab. 5 presents FGBERT’s classification results on NCycDB, suggesting its ability to recognize key features of particular N cycle processes and improve gene family classification by recognizing domains. Although baselines show improved performance on NCycDB compared to PATRIC, due to a larger amount of data per category aiding in discrimination among diverse categories, FGBERT still leads with macro F1 (M.F1) and weighted F1 (W.F1) scores of 99.49% and 99.22%, respectively. However, pre-trained baselines require more time and memory for tokenising large datasets, as analyzed in Sec. 4.3.